# APIGen-MT: Agentic PIpeline for Multi-Turn Data Generation via Simulated Agent-Human Interplay

### v2

@@ Abstract
Training effective AI agents for multi-turn interactions requires high-quality data that captures realistic human-agent dynamics, yet such data is scarce and expensive to collect manually. We introduce APIGen-MT, a two-phase framework that generates verifiable and diverse multi-turn agent data. In the first phase, our agentic pipeline produces detailed task blueprints with ground-truth actions, leveraging a committee of LLM reviewers and iterative feedback loops. These blueprints are then transformed into complete interaction trajectories through simulated human-agent interplay. We train a family of models—the xLAM-2-fc-r series with sizes ranging from 1B to 70B parameters. Our models outperform frontier models such as GPT-4o and Claude 3.5 on $\tau$-bench and BFCL benchmarks, with the smaller models surpassing their larger counterparts, particularly in multi-turn settings, while maintaining superior consistency across multiple trials. Comprehensive experiments demonstrate that our verified blueprint-to-details approach yields high-quality training data, enabling the development of more reliable, efficient, and capable agents. We open-source both the synthetic data collected and the trained xLAM-2-fc-r models to advance research in AI agents.
@@
[Uncaptioned image] Website & Dataset https://apigen-mt.github.io
@@ 1 Introduction
  • We open-source our high-quality synthetic data and trained models, i.e., the xLAM-2-fc-r series, to advance research in AI agent space.
@@ 3.1 Multi-Turn Interaction Problem Formulation
Multi-turn interactions between an AI assistant and a human user present unique challenges that go beyond single-turn exchanges. We formalize this interaction as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(\mathcal{U},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},\mathcal{R})$, where $\mathcal{U}$ represents the instruction space containing possible user intents; $\mathcal{S}$ denotes the state space of the environment and conversation history; $\mathcal{A}=\{tool\_call,response\}$ is the action space available to the assistant; $\mathcal{O}=\mathcal{O}_{E}\cup\mathcal{O}_{H}$ is the observation space comprising observations from the environment ($\mathcal{O}_{E}$) and response from the human ($\mathcal{O}_{H}$); $\mathcal{T}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}\times\mathcal{O}$ is the transition function; and $\mathcal{R}$ is the reward function evaluating interaction success.
The AI assistant must engage in a multi-turn conversation with the human user to incrementally understand their intent $q\in\mathcal{U}$ and solve it through appropriate interactions with the environment while adhering to any domain rules. At turn $t$, the assistant predicts an action $a^{t}\in\mathcal{A}$ based on the interaction history and understanding of $q$ thus far. When $a^{t}$ is a $tool\_call$ compliant with the rules, it triggers a state transition $(s^{t}_{E},tool\_call)\rightarrow(s^{t+1}_{E},o_{E})$, where $o_{E}\in\mathcal{O}_{E}$ is the tool output (typically in structured format like JSON). When $a^{t}$ is a $response$ to the human, it causes a state transition $(s^{t}_{H},response)\rightarrow(s^{t+1}_{H},o_{H})$, where $o_{H}\in\mathcal{O}_{H}$ is the human’s follow-up message. Importantly, the environment state $s^{t+1}_{E}$ remains latent to both the assistant and the human. The interaction completes when the human sends a terminating message or the maximum number of turns is reached. The reward $\mathcal{R}(\Delta\mathcal{S}_{E},a)$ is calculated based on the cumulative state change in the environment $\Delta\mathcal{S}_{E}$ and the sequence of responses $a=\{a_{i}\mid a_{i}\in response\text{ to human}\}$ provided by the assistant throughout the episode. The assistant’s objective is to maximize this reward.
@@ 4.3 Data Collection & Statistics
Data Collection Procedure. We source APIs implemented as Python functions from $\tau$-bench. Among these, we have 15 ‘read’ and 13 ‘write’ APIs across both domains. $\tau$-bench is accompanied with detailed policies and domain rules in two settings - Retail and Airline which we use as guideline policies. We utilize gpt-4o and DeepSeek V3 and R1 models in the task generation, validation and agent-human interplay stages to collect training data. The prompts used in every stage are provided in Appendix B. We set the maximum number of reflection-based feedback turns to 3 for retail and 5 for airline respectively.
@@ 5.2 Experiment Results
Table 1: Performance comparison of different models on BFCL leaderboard (as of date 04/03/2025). The rank is based on the overall accuracy, which is a weighted average of different evaluation categories. “FC" stands for function-calling mode in contrast to using a customized “prompt" to extract the function calls. See the benchmark [45] for details.
@@
Table 2 shows performance under the default naive user simulation setting on $\tau$-bench. Our xLAM-2-70b-fc-r model achieves an overall success rate of 56.2%, significantly outperforming the base Llama 3.1 70B Instruct model (38.2%) and other open-source models like DeepSeek v3 (40.6%). Notably, our model even outperforms proprietary models such as gpt-4o (52.9%) and approaches the performance of more recent models like Claude 3.5 Sonnet (new) (60.1%).
@@
A particularly impressive result is that our smaller models, such as xLAM-2-32b-fc-r and xLAM-2-8b-fc-r, achieve remarkable performance (54.6% and 46.7% respectively), outperforming much larger baseline models. This suggests that our synthetic data generation approach enables efficient knowledge transfer, allowing smaller models to achieve competitive performance with significantly fewer parameters.
@@ 5.4 In-Depth Analysis of Model Behavior
To better understand the behavior of our trained models, we perform an in-depth investigation of the tasks solved by xLAM-2-70b-fc-r and a state-of-the-art model Claude 3.5 Sonnet (new) on $\tau$-bench. We categorize tasks into ‘short’, ‘medium’ and ‘long’ based on the number of turns required by Claude 3.5 to solve each task across a union of 8 trials. This categorization is derived by calculating the 33rd and 66th percentiles of the number of turns.
Figure 7: Performance and efficiency comparisons of xLAM-2-70b-fc-r with frontier models on $\tau$-bench.
From Figure 7 we observe that particularly on the ‘long’ task category, the success rate for xLAM-2-70b-fc-r is much higher than gpt-4o but lags behind Claude. Further, we assess the efficiency of the agent by measuring the number of interactions needed with the simulated user for the agent to fully comprehend the intent and successfully complete the task. The plot reveals that xLAM-2-70b-fc-r is at par with gpt-4o but requires more interactions compared to Claude, which can be attributed to its method of retrieving user details in stages, necessitating more turns. These observations suggest potential areas for improvement in future iterations.

### v4

@@ Abstract
Training effective AI agents for multi-turn interactions requires high-quality data that captures realistic human-agent dynamics, yet such data is scarce and expensive to collect manually. We introduce APIGen-MT, a two-phase framework that generates verifiable and diverse multi-turn agent data. In the first phase, our agentic pipeline produces detailed task blueprints with ground-truth actions, leveraging a committee of LLM reviewers and iterative feedback loops. These blueprints are then transformed into complete interaction trajectories through simulated human-agent interplay. We train a family of models—the xLAM-2-fc-r series with sizes ranging from 1B to 70B parameters. Our models outperform frontier models such as GPT-4o and Claude 3.5 on $\tau$-bench and BFCL benchmarks, with the smaller models surpassing their larger counterparts, particularly in multi-turn settings, while maintaining superior consistency across multiple trials. Comprehensive experiments demonstrate that our verified blueprint-to-details approach yields high-quality training data, enabling the development of more reliable, efficient, and capable agents. We open-source 5K synthetic data trajectories and the trained xLAM-2-fc-r models to advance research in AI agents.
@@
[Uncaptioned image] Dataset https://huggingface.co/Salesforce/APIGen-MT-5k
[Uncaptioned image] Website https://apigen-mt.github.io
@@ 1 Introduction
  • We open-source 5K high-quality synthetic data (APIGen-MT-5k) and trained models, i.e., the xLAM-2-fc-r series, to advance research in AI agent space.
@@ 3.1 Multi-Turn Interaction Problem Formulation
Multi-turn interactions between an AI assistant and a human user present unique challenges that go beyond single-turn exchanges. We formalize this interaction as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(\mathcal{U},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},\mathcal{R})$, where $\mathcal{U}$ represents the instruction space containing possible user intents; $\mathcal{S}$ denotes the state space of the environment and conversation history; $\mathcal{A}=\{tool\_call,response\}$ is the action space available to the assistant; $\mathcal{O}=\mathcal{O}_{E}\cup\mathcal{O}_{H}$ is the observation space comprising observations from the environment ($\mathcal{O}_{E}$) and response from the human ($\mathcal{O}_{H}$); $\mathcal{T}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}\times\mathcal{O}$ is the transition function; and $\mathcal{R}$ is the reward function evaluating interaction success. The AI assistant must engage in a multi-turn conversation with the human user to incrementally understand their intent $q\in\mathcal{U}$ and solve it through appropriate interactions with the environment while adhering to any domain rules. At turn $t$, the assistant predicts an action $a^{t}\in\mathcal{A}$ based on the interaction history and understanding of $q$ thus far. When $a^{t}$ is a $tool\_call$ compliant with the rules, it triggers a state transition $(s^{t}_{E},tool\_call)\rightarrow(s^{t+1}_{E},o_{E})$, where $o_{E}\in\mathcal{O}_{E}$ is the tool output (typically in structured format like JSON). When $a^{t}$ is a $response$ to the human, it causes a state transition $(s^{t}_{H},response)\rightarrow(s^{t+1}_{H},o_{H})$, where $o_{H}\in\mathcal{O}_{H}$ is the human’s follow-up message. Importantly, the environment state $s^{t+1}_{E}$ remains latent to both the assistant and the human. The interaction completes when the human sends a terminating message or the maximum number of turns is reached. The reward $\mathcal{R}(\Delta\mathcal{S}_{E},a)$ is calculated based on the cumulative state change in the environment $\Delta\mathcal{S}_{E}$ and the sequence of responses $a=\{a_{i}\mid a_{i}\in response\text{ to human}\}$ provided by the assistant throughout the episode. The assistant’s objective is to maximize this reward.
@@ 4.3 Data Collection & Statistics
Data Collection Procedure. We source APIs implemented as Python functions from $\tau$-bench. Among these, we have 15 ‘read’ and 13 ‘write’ APIs across both domains. $\tau$-bench is accompanied with detailed policies and domain rules in two settings - Retail and Airline which we use as guideline policies. We utilize gpt-4o and DeepSeek V3 models in the task generation, validation and agent-human interplay stages to collect training data. The prompts used in every stage are provided in Appendix B. We set the maximum number of reflection-based feedback turns to 3 for retail and 5 for airline respectively.
@@ 5.2 Experiment Results
Table 1: Performance of different models on BFCL leaderboard (as of date 04/03/2025). The rank is based on the overall accuracy, which is a weighted average of different evaluation categories. “FC" stands for function-calling mode in contrast to using a customized “prompt" to extract the function calls. See the benchmark [45] for details.
@@
Table 2 presents results under the default naive user setting on $\tau$-bench. Our xLAM-2-70b-fc-r model achieves a 56.2% success rate, outperforming Llama 3.1 70B Instruct (38.2%), DeepSeek v3 (40.6%), and even proprietary models like GPT-4o (52.9%), while approaching more recent models like Claude 3.5 Sonnet (60.1%). Notably, our smaller variants like xLAM-2-32b-fc-r (54.6%) and xLAM-2-8b-fc-r (46.7%) surpass larger baselines, demonstrating that our synthetic data approach enables efficient knowledge transfer and strong performance with fewer parameters.
@@ 5.4 In-Depth Analysis of Model Behavior
Figure 7: Performance/efficiency comparisons of xLAM-2-70b-fc-r with frontier models on $\tau$-bench.
To better understand the behavior of our trained models, we perform an in-depth investigation of the tasks solved by xLAM-2-70b-fc-r and a state-of-the-art model Claude 3.5 Sonnet (new) on $\tau$-bench. We categorize tasks into ‘short’, ‘medium’ and ‘long’ based on the number of turns required by Claude 3.5 to solve each task across a union of 8 trials. This categorization is derived by calculating the 33rd and 66th percentiles of the number of turns. From Figure 7 we observe that particularly on the ‘long’ task category, the success rate for xLAM-2-70b-fc-r is much higher than gpt-4o but lags behind Claude. Further, we assess the efficiency of the agent by measuring the number of interactions needed with the simulated user for the agent to fully comprehend the intent and successfully complete the task. The plot reveals that xLAM-2-70b-fc-r is at par with gpt-4o but requires more interactions compared to Claude, which can be attributed to its method of retrieving user details in stages, necessitating more turns. These observations suggest potential areas for improvement in future iterations.